# Online Hypergraph Matching

### v2

@@ Abstract
The online matching problem was introduced by Karp, Vazirani and Vazirani (STOC 1990) on bipartite graphs with vertex arrivals. It is well-known that the optimal competitive ratio is $1-1/e$ for both integral and fractional versions of the problem. Since then, there has been considerable effort to find optimal competitive ratios for other related settings.
@@ 1 Introduction
Online matching is a classic problem in the field of online algorithms. It was first introduced in the seminal work of Karp, Vazirani and Vazirani [24], who considered the bipartite version with one-sided vertex arrivals. In this setting, we are given a bipartite graph where vertices on one side are known in advance (offline), and vertices on the other side arrive sequentially (online). When an online vertex arrives, it reveals its incident edges, at which point the algorithm must decide how to match it (or not) irrevocably. The goal is to maximize the cardinality of the resulting matching. Karp et al. [24] gave an elegant randomized algorithm Ranking, which achieves the optimal competitive ratio of $1-1/e$.
In certain applications, each offline vertex may be matched more than once. Examples include matching online jobs to servers, or matching online impressions to advertisers. This is the online $b$-matching model of Kalyanasundaram and Pruhs [22], where $b\geq 1$ is the maximum number of times an offline vertex can be matched. As $b$ and the number of online vertices tend to infinity, it in turn captures the fractional relaxation of the Karp et al. [24] model. This means that the algorithm is allowed to match an online node fractionally to multiple neighbours, as long as the total load on every vertex does not exceed 1. For this problem, it is known that the deterministic algorithm Balance (or Water-Filling) achieves the optimal competitive ratio of $1-1/e$.
@@ 1.1 Online hypergraph matching
The online bipartite matching problem can be naturally generalized to hypergraphs as follows. For $k\geq 2$, let $\mathcal{H}=(V,W,H)$ be a $k$-uniform hypergraph with offline vertices $V$, online vertices $W$ and hyperedges $H$. Every hyperedge $h\in H$ contains $k-1$ elements from $V$ and 1 element from $W$. Just like before, the online vertices arrive sequentially with their incident hyperedges, and the goal is to select a large matching, i.e., a set of disjoint hyperedges. The greedy algorithm is $1/k$-competitive. On the other hand, no integral algorithm can be $2/k$-competitive  [^1] In [35], it is shown that no algorithm can be $(2+f(k))/k$-competitive for some positive function $f$ with $f(k)=o(1)$. In Appendix C, we give a simple construction showing that no integral algorithm can be $2/k$-competitive..
@@
The aforementioned results show that asymptotically, both integral and fractional versions of the online matching problem on $k$-uniform hypergraphs are essentially settled (up to constant factors). However, our understanding of the problem for small values of $k$ (other than $k=2$) remains poor. Many applications of online hypergraph matching in practice have small values of $k$. For instance, in ride-sharing and on-demand delivery services [32], $k-1$ represents the capacity of service vehicles, which is often small. Another example is network revenue management problems [27]. In this setting, given a collection of limited resources, a sequence of product requests arrive over time. When a product request arrives, we have to decide whether to accept it irrevocably. Accepting a product request generates profit, but also consumes a certain amount of each resource. The goal is to devise a policy which maximizes profit. In this context, $k-1$ represents the maximum number of resources used by a product. As Ma et al. [27] noted, many of these problems have small values of $k$. In airlines, for example, $k-1$ corresponds to the maximum number of flight legs included in an itinerary, which usually does not exceed two or three.
@@ 1.2 Our contributions
For the online fractional matching problem on $3$-uniform hypergraphs, there is a deterministic $(e-1)/(e+1)$-competitive algorithm. Furthermore, every algorithm is at most $(e-1)/(e+1)$-competitive.
@@
Note that in Theorem 1.2, the first term is at most the second term if and only if $d\leq k-1$. Moreover, Random is at least as good as the greedy algorithm, since the latter is $1/k$-competitive. For $3$-uniform hypergraphs, the bound becomes $1/2$ for $d\leq 2$ and $1/(3-2/d)$ otherwise, thus interpolating between $1/3$ and $1/2$. Note that for $d\leq 2$, the bound is optimal, since the online matching problem on graphs under edge arrivals is a special case of this setting (with $k=3,d=1$), for which an upper bound of $1/2$ is known even against fractional algorithms on bipartite graphs [15].
@@ 1.3 Related work
Since the online matching problem was introduced in [24], it has garnered a lot of interest, leading to extensive follow-up work. We refer the reader to the excellent survey by Mehta [30] for navigating this rich literature. The original analysis of Ranking [24] was simplified in a series of papers [3, 10, 16, 12]. Many variants of the problem have been studied, such as the online $b$-matching problem [22], and its extension to the AdWords problem [4, 9, 21, 31]. Weighted generalizations have been considered, e.g., vertex weights [1, 19] and edge weights [14]. Weakening the adversary by requiring that online nodes arrive in a random order has also been of interest [23, 28, 25]. Another line of research explored more general arrival models such as two-sided vertex arrival [36], general vertex arrival [15], edge arrival [6, 15], and general vertex arrival with departure times [18, 20, 2].
@@
Hypergraph matching on $k$-uniform hypergraphs is a well-studied problem in the offline setting. It is NP-hard to approximate within a factor of $\Omega(\log(k)/k)$ [17].Moreover, the factor between the optimal solution and the optimal value of the natural LP relaxation is at least $1/(k-1+1/k)$ [7].
A special case that has also been studied is the restriction to $k$-partite graphs, where the vertices are partitioned into $k$ sets and every hyperedge contains exactly one vertex from each set. This setting is called $k$-dimensional matching, and the optimal solution is known to be at least $1/(k-1)$ times the optimal value of the standard LP relaxation [7]. For $k=3$, the best known polynomial time approximation algorithm gives a $(3/4-\varepsilon)$-approximation [8].
@@ 2 Preliminaries
The online matching problem on $k$-uniform hypergraphs under vertex arrivals is defined as follows. An instance consists of a $k$-uniform hypergraph $\mathcal{H}=(V,W,H)$, where $V$ is the set of offline nodes and $W=(w_{1},w_{2},\ldots)$ is the sequence of online nodes. The ordering of $W$ corresponds to the arrival order of the online nodes. Every hyperedge $h\in H$ has exactly one node in $W$ and $k-1$ nodes in $V$. When an online node $w\in W$ arrives, its incident hyperedges $\delta(w)$ are revealed. A fractional algorithm is allowed to irrevocably increase $x_{h}$ for every $h\in\delta(w)$, whereas an integral algorithm is allowed to irrevocably pick one of these hyperedges, i.e., setting $x_{h}=1$ for some $h\in\delta(w)$.
@@ 3 Optimal fractional algorithm for 3-uniform hypergraphs
In this section, we present a primal-dual algorithm for the online fractional matching problem on 3-uniform hypergraphs under vertex arrivals. This algorithm will turn out to be optimal with a tight competitive ratio of $(e-1)/(e+1)\approx 0.4621$. We define the following distribution function $f:[0,1]\to[0,1]$:
$f(x):=\frac{e^{x}}{e+1}.$ (2)
@@
$\phi(h):=f(x(\delta(u)))+f(x(\delta(v))).$ (3)
Figure 1 shows the possible values of $x(\delta(u))$ and $x(\delta(v))$ such that $\phi(h)\leq 1$. We now present the algorithm.
@@
$\mathbf{Input:}$ $3$-uniform hypergraph $\mathcal{H}=(V,W,H)$ with online nodes $W$.
@@
  increase $x_{h}$ for every $h=\{u,v,w\}\in\arg\min_{h\in\delta(w)}\{\phi(h)\}$ at rate $1$
@@
Algorithm 3.1 is $(e-1)/(e+1)$-competitive for the online fractional matching problem on $3$-uniform hypergraphs.
@@
$\mathcal{V}(\mathcal{A})=\sum_{h\in H}x_{h}$ $=\sum_{v\in V\cup W}y_{v}\qquad\text{and}$ (4)
$\sum_{v\in h}y_{v}$ $\geq\rho\qquad\forall h\in H.$ (5)
@@ 4 Tight upper bound for 3-uniform hypergraphs
We now prove the second part of Theorem 1.1, i.e., every algorithm is at most $(e-1)/(e+1)$-competitive for the online fractional matching problem on 3-uniform hypergraphs under vertex arrivals.
@@ 4.1 Overview of the construction
We construct an adversarial instance that is adaptive to the behaviour of the algorithm. The main idea is to combine the vertex-arrival instance of Karp et al. [24] and the edge-arrival instance of Gamlath et al. [15] for bipartite graphs.
We start by giving a high-level overview of the construction. The offline vertices of the hypergraph are partitioned into $m$ sets $C_{1},\dots,C_{m}$, which we call components. Each component will induce a bipartite graph with bipartition $C_{i}=U_{i}\cup V_{i}$, where $|U_{i}|=|V_{i}|=T$.
@@
    On each component $C_{i}$, the matching $\mathcal{M}_{i}^{(t)}$ is constructed based on the behaviour of the algorithm in phase $t-1$. It draws inspiration from the edge-arrival instance in [15], together with the function $f(x)=e^{x}/(e+1)$ defined in (2). The exact construction is described in Section 4.3 and illustrated in Figures 2 and 3.
@@ 4.2 Assumptions on the algorithm
    The algorithm is $\varepsilon$-threshold respecting on all online nodes in the first $T-1$ phases for some arbitrarily small $\varepsilon>0$.
@@ 4.3 Constructing the matching $\mathcal{M}^{(t)}$
In this section, we construct the matching $\mathcal{M}_{i}^{(t)}$ for every component $i\in[m]$ and phase $t\in[T]$. We will only describe the matchings for a single component $C_{i}$, i.e., $\mathcal{M}_{i}^{(1)},\mathcal{M}_{i}^{(2)},\dots,\mathcal{M}_{i}^{(T)}$, because the same construction applies to other components. Intuitively, the value obtained by the algorithm in the first $T-1$ phases is already limited by Assumption 1. So, the goal of this construction is to prevent the algorithm from gaining too much value in the last phase $T$. In particular, we will show that it can only obtain $O(\sqrt{T})+\varepsilon\;O(T^{2})$ in the last phase on every component $C_{i}$.
Figure 2: The partial matchings $\mathcal{M}_{i}^{(t)}$ if the fractional algorithm ensures that every edge reaches the threshold at the end of every phase, meaning that $f(\ell_{u}^{(t)})+f(\ell_{v}^{(t)})\geq 1$ for every $(u,v)\in\mathcal{M}_{i}^{(t)}.$ The maximum matching at the end of phase $5$ has size five and consists of $\mathcal{M}_{i}^{(5)}$.
@@
    where we define $\sigma_{t}(r_{t}+1):=t+1$ for convenience. In particular, note that $t+1\in U_{i}$ and $t+1\in V_{i}$ are two fresh nodes with zero load, which are always part of the matching $\mathcal{M}_{i}^{(t+1)}$, but not part of any matching from a previous phase. Clearly, the invariant is maintained. Figures 2 and 3 illustrate the construction.
Figure 3: In this example, the algorithm does not increase the edge $(1,3)$, and thus by symmetry the edge $(3,1)$, up to the threshold during phase $t=3$. Hence, $f(\ell_{1}^{(3)})+f(\ell_{3}^{(3)})<1$ and nodes $1$ and $3$ become inactive from that point on. The maximum matching at the end of phase $5$ still has size five and consists of $\mathcal{M}_{i}^{(5)}$, in addition to the two edges $(1,3)$ and $(3,1)$ that are below the threshold.

### v3

# Sander Borst Danish Kashaev Zhuan Khye Koh

@@ Abstract
The online matching problem was introduced by Karp, Vazirani and Vazirani (STOC 1990) on bipartite graphs with vertex arrivals. It is well-known that the optimal competitive ratio is 11/e1-1/e for both integral and fractional versions of the problem. Since then, there has been considerable effort to find optimal competitive ratios for other related settings.
@@ 1 Introduction
Online matching is a classic problem in the field of online algorithms. It was first introduced in the seminal work of Karp, Vazirani and Vazirani [23], who considered the bipartite version with one-sided vertex arrivals. In this setting, we are given a bipartite graph where vertices on one side are known in advance (offline), and vertices on the other side arrive sequentially (online). When an online vertex arrives, it reveals its incident edges, at which point the algorithm must decide how to match it (or not) irrevocably. The goal is to maximize the cardinality of the resulting matching. Karp et al. [23] gave an elegant randomized algorithm Ranking, which achieves the optimal competitive ratio of 11/e1-1/e.
In certain applications, each offline vertex may be matched more than once. Examples include matching online jobs to servers, or matching online impressions to advertisers. This is the online $b$-matching model of Kalyanasundaram and Pruhs [21], where $b\geq 1$ is the maximum number of times an offline vertex can be matched. As $b$ and the number of online vertices tend to infinity, it in turn captures the fractional relaxation of the Karp et al. [23] model. This means that the algorithm is allowed to match an online node fractionally to multiple neighbours, as long as the total load on every vertex does not exceed 1. For this problem, it is known that the deterministic algorithm Balance (or Water-Filling) achieves the optimal competitive ratio of 11/e1-1/e.
@@ 1.1 Online hypergraph matching
The online bipartite matching problem can be naturally generalized to hypergraphs as follows. For $k\geq 2$, let $\mathcal{H}=(V,W,H)$ be a $k$-uniform hypergraph with offline vertices $V$, online vertices $W$ and hyperedges $H$. Every hyperedge $h\in H$ contains k1k-1 elements from $V$ and 1 element from $W$. Just like before, the online vertices arrive sequentially with their incident hyperedges, and the goal is to select a large matching, i.e., a set of disjoint hyperedges. The greedy algorithm is $1/k$-competitive. On the other hand, no integral algorithm can be $2/k$-competitive  [^1] In [35], it is shown that no algorithm can be $(2+f(k))/k$-competitive for some positive function $f$ with $f(k)=o(1)$. In Appendix D, we give a simple construction showing that no integral algorithm can be $2/k$-competitive..
@@
The aforementioned results show that asymptotically, both integral and fractional versions of the online matching problem on $k$-uniform hypergraphs are essentially settled (up to constant factors). However, our understanding of the problem for small values of $k$ (other than $k=2$) remains poor. Many applications of online hypergraph matching in practice have small values of $k$. For instance, in ride-sharing and on-demand delivery services [32], k1k-1 represents the capacity of service vehicles, which is often small. Another example is network revenue management problems [27]. In this setting, given a collection of limited resources, a sequence of product requests arrive over time. When a product request arrives, we have to decide whether to accept it irrevocably. Accepting a product request generates profit, but also consumes a certain amount of each resource. The goal is to devise a policy which maximizes profit. In this context, k1k-1 represents the maximum number of resources used by a product. As Ma et al. [27] noted, many of these problems have small values of $k$. In airlines, for example, k1k-1 corresponds to the maximum number of flight legs included in an itinerary, which usually does not exceed two or three.
@@ 1.2 Our contributions
For the online fractional matching problem on $3$-uniform hypergraphs, there is a deterministic (e1)/(e+1)(e-1)/(e+1)-competitive algorithm. Furthermore, every algorithm is at most (e1)/(e+1)(e-1)/(e+1)-competitive.
@@
Note that in Theorem 1.2, the first term is at most the second term if and only if $d\leq k-1$. Moreover, Random is at least as good as the greedy algorithm, since the latter is $1/k$-competitive. For $3$-uniform hypergraphs, the bound becomes $1/2$ for $d\leq 2$ and 1/(32/d)1/(3-2/d) otherwise, thus interpolating between $1/3$ and $1/2$. Note that for $d\leq 2$, the bound is optimal, since the online matching problem on graphs under edge arrivals is a special case of this setting (with $k=3,d=1$), for which an upper bound of $1/2$ is known even against fractional algorithms on bipartite graphs [15].
@@ 1.3 Related work
Since the online matching problem was introduced in [23], it has garnered significant interest, leading to extensive follow-up work. We refer the reader to the excellent survey by Mehta [31] for navigating this rich literature. The original analysis of Ranking [23] was simplified in a series of papers [3, 9, 16, 12]. Many variants of the problem have been studied, such as the online $b$-matching problem [21], and its extension to the AdWords problem [4, 10, 20, 30]. Weighted generalizations have been considered, e.g., vertex weights [1, 18] and edge weights [14]. Weakening the adversary by requiring that online nodes arrive in a random order has also been of interest [22, 28, 24]. Another line of research explored more general arrival models such as two-sided vertex arrival [36], general vertex arrival [15], edge arrival [6, 15], and general vertex arrival with departure times [17, 19, 2].
@@
Hypergraph matching on $k$-uniform hypergraphs is a well-studied problem in the offline setting. It is known to be impossible to approximate within a factor of $\Omega(1/k)$ in polynomial time, unless $\mathbf{NP}\subseteq\mathbf{BPP}$ [26]. Moreover, the factor between the optimal solution and the optimal value of the natural LP relaxation is at least 1/(k1+1/k)1/(k-1+1/k) [7].
A special case that has also been studied is the restriction to $k$-partite graphs, where the vertices are partitioned into $k$ sets and every hyperedge contains exactly one vertex from each set. This setting is called $k$-dimensional matching, and the optimal solution is known to be at least 1/(k1)1/(k-1) times the optimal value of the standard LP relaxation [7]. For $k=3$, the best known polynomial time approximation algorithm gives a $(3/4-\varepsilon)$-approximation [8].
@@ 2 Preliminaries
The online matching problem on $k$-uniform hypergraphs under vertex arrivals is defined as follows. An instance consists of a $k$-uniform hypergraph $\mathcal{H}=(V,W,H)$, where $V$ is the set of offline nodes and $W=(w_{1},w_{2},\ldots)$ is the sequence of online nodes. The ordering of $W$ corresponds to the arrival order of the online nodes. Every hyperedge $h\in H$ has exactly one node in $W$ and k1k-1 nodes in $V$. When an online node $w\in W$ arrives, its incident hyperedges $\delta(w)$ are revealed. A fractional algorithm is allowed to irrevocably increase $x_{h}$ for every $h\in\delta(w)$, whereas an integral algorithm is allowed to irrevocably pick one of these hyperedges, i.e., setting $x_{h}=1$ for some $h\in\delta(w)$.
@@ 3 Optimal fractional algorithm for 3-uniform hypergraphs
In this section, we present a primal-dual algorithm for the online fractional matching problem on 3-uniform hypergraphs under vertex arrivals. This algorithm will turn out to be optimal with a tight competitive ratio of $(e-1)/(e+1)\approx 0.4621$. We define the following function $f:[0,1]\to[0,1]$:
[…]
@@
[…]
Figure 3 shows the possible values of $x(\delta(u))$ and $x(\delta(v))$ such that $\phi(h)\leq 1$. We now present the algorithm. The dual updates are shaded in gray since they are only needed for the analysis.
@@
$\mathbf{Input:}$ 33-uniform hypergraph $\mathcal{H}=(V,W,H)$ with online nodes $W$.
@@
 increase $x_{h}$ for every $h=\{u,v,w\}\in\arg\min_{h\in\delta(w)}\{\phi(h)\}$ at rate 11
@@
Algorithm 3.1 is (e1)/(e+1)(e-1)/(e+1)-competitive for the online fractional matching problem on $3$-uniform hypergraphs.
@@
[…]
@@
We note that the above algorithm and its analysis are very similar to those of [5]. Here we optimized the choice of $f$ for the case of 3-uniform hypergraphs. By choosing $f(x):=(k\log(k))^{x-1}$ and setting $\phi(h):=\sum_{u\in h}f(x(\delta(u)))$ for every $k$-uniform hyperedge $h$, it can be shown that the natural generalization of Algorithm 3.1 to $k$-uniform hypergraphs is $\Omega(1/\log(k))$-competitive [35].
@@ 4 Tight upper bound for 3-uniform hypergraphs
We now prove the second part of Theorem 1.1, i.e., every algorithm is at most (e1)/(e+1)(e-1)/(e+1)-competitive for the online fractional matching problem on 3-uniform hypergraphs under vertex arrivals.
@@ 4.1 Overview of the construction
We construct an adversarial instance that is adaptive to the behaviour of the algorithm. The main idea is to combine the vertex-arrival instance of Karp et al. [23] and the edge-arrival instance of Gamlath et al. [15] for bipartite graphs. We first briefly describe these two instances, and then give a high-level view of our construction.
Bipartite hardness under vertex-arrivals.
An illustration of this construction is given in Figure 1. This instance consists of a bipartite graph, where one side of the bipartition has size $n\in\mathbb{N}$ and is known upfront. The other side of the bipartition arrives online, with $n$ online nodes arriving in total. The $k$th online node has degree nk+1n-k+1, with the edges incident to it connecting to the nk+1n-k+1 offline nodes with the highest fractional degree at that point in time.
Note that, after all online nodes have arrived, the instance contains a perfect matching of size $n$. It can however be shown that no fractional algorithm can achieve a competitive ratio better than $1-1/e\approx 0.63$ for this instance, due to the uncertainty about the edge adjacent to each online node in the perfect matching, since it is a decision made online by the adversary. This construction can easily be adapted to our model on $3$-uniform hypergraphs under vertex arrivals, by replacing each offline node by a pair of two identical offline nodes, and replacing each edge of the graph by a hyperedge of size three.
Figure 1: The hard instance for bipartite graphs under vertex-arrivals. The $k$th online node connects to the nk+1n-k+1 offline nodes with the highest fractional degree at that point in time. In this figure, the offline vertices on the left side of the bipartition are drawn in decreasing order of their fractional degrees. The perfect matching consists of all the horizontal edges.
Figure 2: The hard instance for bipartite graphs under edge-arrivals. The adversary can terminate the instance at any phase if the algorithm does not generate enough value. The optimal matching consists of all the edges which arrived in the last phase.
Bipartite hardness under edge-arrivals.
An illustration of this construction is given in Figure 2. Each side of the bipartition has cardinality $T$ and the nodes on both sides are denoted by $\{1,\dots,T\}$. This instance now consists of $T^{\prime}\leq T$ phases, where $T^{\prime}$ is unknown a priori and determined online by the adversary given the choices of the algorithm. At each phase $t\in[T^{\prime}]$, the following $t$ new edges, which form a matching, are revealed:
$\Big\{(i,t-i+1):i\in\{1,\dots,t\}\Big\}.$
Observe that these become the optimal matching at that point in time. If the value generated by the algorithm exceeds $\alpha t$ at the end of phase $t$ for some $\alpha\in[0,1]$, the instance continues to the next phase. Otherwise, the instance terminates. It is shown in [15] that no fractional algorithm can achieve a competitive ratio better than $1/2$ for this instance. The hardness is due to the uncertainty about the time horizon: the adversary can always make a new set of edges arrive if the algorithm gains enough value in each phase. This construction can easily be adapted to our vertex-arrival model on $3$-uniform hypergraphs, by appending an online node with degree $1$ to each edge of the bipartite graph.
Our construction.
We now give a high-level overview of our construction, which combines both hard instances described above. The offline vertices of the hypergraph are partitioned into $m$ sets $C_{1},\dots,C_{m}$, which we call components. Each component will induce a bipartite graph with bipartition $C_{i}=U_{i}\cup V_{i}$, where $U_{i}$ and $V_{i}$ are both ordered vertex sets of size $T\in\mathbb{N}$.
@@
    On each component $C_{i}$, the matching $\mathcal{M}_{i}^{(t)}$ is constructed based on the behaviour of the algorithm in phase t1t-1. It draws inspiration from the edge-arrival instance in [15], together with the function $f(x)=e^{x}/(e+1)$ defined in (2). The exact construction is described in Section 4.3 and illustrated in Figure 4.
@@ 4.2 Assumptions on the algorithm
    The algorithm is $\varepsilon$-threshold respecting on all online nodes in the first T1T-1 phases for some arbitrarily small $\varepsilon>0$.
@@ 4.3 Constructing the matching $\mathcal{M}^{(t)}$
In this section, we construct the matching $\mathcal{M}_{i}^{(t)}$ for every component $i\in[m]$ and phase $t\in[T]$. We will only describe the matchings for a single component $C_{i}$, i.e., $\mathcal{M}_{i}^{(1)},\mathcal{M}_{i}^{(2)},\dots,\mathcal{M}_{i}^{(T)}$, because the same construction applies to other components. Intuitively, the value obtained by the algorithm in the first T1T-1 phases is already limited by Assumption 1. So, the goal of this construction is to prevent the algorithm from gaining too much value in the last phase $T$. In particular, we will show that it can only obtain $O(\sqrt{T})+\varepsilon\;O(T^{2})$ in the last phase on every component $C_{i}$.
@@
    where we define $\sigma_{t}(r_{t}+1):=t+1$ for convenience. In particular, note that $t+1\in U_{i}$ and $t+1\in V_{i}$ are two fresh nodes with zero load, which are always part of the matching $\mathcal{M}_{i}^{(t+1)}$, but not part of any matching from a previous phase. Clearly, the invariant is maintained. Figure 4 illustrates the construction.
Figure 4: In this example, the algorithm does not increase the edge $(1,3)$, and thus by symmetry the edge $(3,1)$, up to the threshold during phase $t=3$. Hence, $f(\ell_{1}^{(3)})+f(\ell_{3}^{(3)})<1$ and nodes $1$ and 33 become inactive from that point on. The maximum matching at the end of phase 55 still has size five and consists of $\mathcal{M}_{i}^{(5)}$, in addition to the two edges $(1,3)$ and $(3,1)$ that are below the threshold.